OpenStudy (anonymous):

Solve the equation. x^3 − 7x^2 + 12x = 0 X= x= x=

OpenStudy (anonymous):

x(x-3)(x-4)

OpenStudy (anonymous):

x^3 − 7x^2 + 12x = 0 x(x^2-7x+12)=0 factor gcf x(x-3)(x-4)=0 factor trinomial (-3*-4=12; -3-4=-7) x=0 or x-3=0 or x-4=0 zero-factor theorem x=0 or x=3 or x=4 solve {0, 3, 4} solution set